OpenStudy (anonymous):

An integer N is two thirds of another and their difference is 10. Find the two Integers. (answer needed asap plz)

OpenStudy (anonymous):

30,20

OpenStudy (anonymous):

seriously?

OpenStudy (anonymous):

how du do tht?

OpenStudy (anonymous):

Let N be the first number and let x be the second number. Then, N=(2/3)*x Also x-N=10 so x=10+N N=2/3*(x)=2/3*(N+10) N=(2/3)*N+20/3 N/3=20/3 N=20 x=10+N=30

OpenStudy (anonymous):

ok will check on it thank you

OpenStudy (anonymous):

welcome